A body of mass 2500g moves through acceleration of 8m/s^2 and 2.2 hours. Determine the momentum?
1
Expert's answer
2020-02-10T09:07:34-0500

The momentum:


"p=mv=mat"

"p=(2.5)(8)(2.2\cdot 3600)=158400\frac{kg\cdot m}{s}"